# Tidally Heated Sub-Neptunes, Refined Planetary Compositions, and Confirmation of a Third Planet in the TOI-1266 System

### v1

@@ 6.2 Joint TTV + RV Modeling
Next, we performed a joint fit of the TTV observations from Table 4 and the RV observations from C24 using TTVFast and emcee, with the same 22 free parameters as described in §6.1. We initialized the 15 planetary parameters close to the location of the best fit solution from the 3-planet TTV fit, and the 7 RV parameters close to the best fit solution from the 3-planet RV fit. We sampled the posterior distribution with DEMCMC for steps, which corresponded to 10 autorrelation lengths for all parameters after discarding the first steps as burn-in. We summarize the prior and posterior distributions for all parameters in Table 3.
@@
Measured planetary parameters, TTV only fit | ||||
@@ 7.4 Planetary Eccentricities and Tides
Next, we assess the potential importance of tidal heating caused by the nonzero eccentricity for the overall energy budgets of planets b and c. For TOI-1266 b, we assume a lower bound of 0.0001 on the eccentricity from our simulations of tidal damping and otherwise draw eccentricity values directly from our TTV+RV posterior distribution. For TOI-1266 c, we also assume a lower bound of 0.0001 from our tidal damping simulations. But since the TTV+RV posterior distribution does not indicate any preference for a nonzero eccentricity of TOI-1266 c, we instead impose an upper bound of 0.001 representing the largest values from our tidal damping simulations. We then calculated the range of tidal heat fluxes from Equation 4 of Jackson et al. (2008) for both planets and plot the resulting distribution in Figure 14. We find that for and combinations that yield a tidal circularization timescale larger than 2 Gyr and are therefore consistent with the larger eccentricity values preferred for planet b, the peak of the distribution for tidal heat flux per unit mass is three orders of magnitude larger than that of Io ( W/kg). This suggests that tidal heating is likely to be important for interpreting the internal and atmospheric composition of TOI-1266 b, with potential implications for long-term atmospheric evolution as well.
@@ 8.1 Bulk Density Constraints
Our updated planet masses and radii do not change the fundamental interpretations of the planet compositions based on the bulk density constraints from C24. The bulk density of TOI-1266 b is too low to be explained solely by rock-water mixtures given the expected upper limit on water mass fraction (WMF) of 50% from planet formation models. This implies that TOI-1266 b’s atmosphere must contain some hydrogen and helium. TOI-1266 c’s smaller radius and larger mass relative to the values reported in C24 now make it easier to match with water-rich scenarios, although it is also consistent with a modest hydrogen-rich envelope (see §8.2). Our updated mass constraint for TOI-1266 c increases the bulk density from 1.6 0.6 g/cm3 to 2.0 0.6 g/cm3, reducing the water-to-rock ratio required to explain the observed mass and radius to 30%. Part of the shift to smaller radii and larger masses for TOI-1266 b and c is also explained by our choice to adopt the stellar parameters from S20, which are based on effective temperature and surface gravity constraints from spectroscopic observations, while C24 adopted stellar parameters from the TIC v8.2 catalog. S20’s stellar mass is larger and their stellar radius is smaller than the values reported in C24.
@@ 8.3 Possible Explanations for the Inverted Architecture of TOI-1266
This picture is additionally complicated by the discovery that TOI-1266 b has a significant tidal heat flux. There are very few compact multi-planet systems with eccentrities as large as our preferred value for TOI-1266 b (e.g., Van Eylen et al., 2019; He et al., 2020). TOI-1266 b’s large tidal heat flux means that its atmosphere may be less massive and/or more metal-rich than implied by the models described in §8, and might help to reduce the disparity in inferred envelope masses relative to TOI-1266 c.
@@ 9.1 TTV Follow-up
There is still much to be gained by additional TTV monitoring of the TOI-1266 system. As shown in Figure 7, we have yet to observe a full TTV super-period for TOI-1266 b. This super-period is caused by its proximity to the 3:1 resonance with TOI-1266 d and has a predicted duration of 36 years (Deck & Agol, 2016), while the current observational baseline is only 4 years. This suggests that future TESS and ground-based transit observations over the next decade could significantly improve on current TTV-based mass and eccentricity constraints. To that end, we provide a list of predicted transit times for all three planets from our dynamical fit in Table 5 in the Appendix.
For planets with circular orbits, there is a well-known degeneracy between the planet masses and eccentricities when fitting the observed TTV signal (Lithwick et al., 2012). The nonzero orbital eccentricities of planets b and/or d create a short-period chopping signature in the measured TTVs for planet b, allowing us to break this degeneracy in our TTV fits.
@@ 10 Summary and Conclusions
Our updated planetary interior and composition modeling confirms that TOI-1266 b requires a hydrogen-rich envelope to explain its observed mass and radius. If we assume that this planet has an Earth-like rocky core and a solar composition envelope, this corresponds to a H/He atmospheric mass fraction of %. TOI-1266 c has a modestly higher bulk density ( = 2.24 g/cc, = 1.54 g/cc) and can therefore be matched with either a small hydrogen-rich envelope with an atmospheric mass fraction ( at 95% confidence), or a water-rich envelope with a mass fraction of % from interior structure models, though we caution that much of this range is unsupported by formation models which predict maximum water mass fractions for terrestrial planets . This picture is additionally complicated by our discovery that planet b has a significantly nonzero orbital eccentricity, corresponding to a large tidal heat flux. We simulate the long-term evolution of the system including the effects of tides and find that TOI-1266 b must have relatively weak tidal dissipation (/) in order to maintain this eccentricity over Gyr timescales. In the future, transmission spectroscopy observations of TOI-1266 b with JWST could directly constrain its internal heat flux, while equivalent observations of TOI 1266 c are needed in order to determine the bulk water content of its envelope. Both planets are among the most favorable temperate sub-Neptune candidates for transmission spectroscopy with JWST, and a detection of atmospheric absorption for either planet would significantly improve our understanding of the TOI-1266 system.
@@ Appendix A Transit Times and Posterior Probability Distributions
TOI-1266 b | TESS PM | 22 | 1930.692530 | 0.001270000 |
TOI-1266 b | TESS PM | 23 | 1941.590140 | 0.001480000 |
TOI-1266 b | TESS PM | 25 | 1963.377320 | 0.002670000 |
TOI-1266 b | TESS EM | 57 | 2312.013329 | 0.000771316 |
TOI-1266 b | TESS EM | 58 | 2322.907835 | 0.000357439 |
@@
TOI-1266 b | TESS SEM | 86 | 2627.962889 | 0.001367946 |
TOI-1266 b | TESS SEM | 87 | 2638.857341 | 0.000514139 |
TOI-1266 b | TESS SEM | 87 | 2638.859811 | 0.001039443 |
TOI-1266 b | TESS SEM | 88 | 2649.750889 | 0.001525124 |
TOI-1266 b | WIRC | 89 | 2660.645312 | 0.001736380 |
@@
TOI-1266 b | MuSCAT | 152 | 3347.007129 | 0.002401229 |
TOI-1266 b | MuSCAT3 | 154 | 3368.796296 | 0.001593378 |
TOI-1266 b | MuSCAT3 | 155 | 3379.692673 | 0.002864712 |
TOI-1266 b | MuSCAT3 | 156 | 3390.588036 | 0.002639571 |
TOI-1266 c | Demory | 0 | 1689.960319 | 0.042524505 |
@@
TOI-1266 c | WIRC | 51 | 2648.845373 | 0.003014198 |
TOI-1266 c | MuSCAT3 | 56 | 2742.873075 | 0.015162918 |
TOI-1266 c | MuSCAT | 89 | 3363.304827 | 0.003415930 |
TOI-1266 c | Sinistro | 90 | 3382.112135 | 0.002607625 |
@@
Figures 17 and 18 show corner plots of the posterior distributions from our joint fit to the TTV and RV data with a 3-planet model. The fit parameters in our fits were planet-to-star mass ratio and , , but we have converted these distributions into units of Earth masses and eccentricity for ease of reference in Figure 17.

### v3

@@ 6.2 Joint TTV + RV Modeling
Next, we performed a joint fit of the TTV observations from Table 4 and the RV observations from C24 using TTVFast and emcee, with the same 22 free parameters as described in §6.1. We initialized the 15 planetary parameters close to the location of the best fit solution from the 3-planet TTV fit, and the 7 RV parameters close to the best fit solution from the 3-planet RV fit. We sampled the posterior distribution with DEMCMC for steps, which corresponded to 10 autocorrelation lengths for all parameters after discarding the first steps as burn-in. We summarize the prior and posterior distributions for all parameters in Table 3.
@@
Measured planetary parameters, TTV only fit (3-planet case) | ||||
@@ 7.4 Planetary Eccentricities and Tides
Next, we assess the potential importance of tidal heating caused by the nonzero eccentricity for the overall energy budgets of planets b and c. For TOI-1266 b, we assume a lower bound of 0.0001 on the eccentricity from our simulations of tidal damping and otherwise draw eccentricity values directly from our TTV+RV posterior distribution. For TOI-1266 c, we also assume a lower bound of 0.0001 from our tidal damping simulations. But since the TTV+RV posterior distribution does not indicate any preference for a nonzero eccentricity of TOI-1266 c, we instead impose an upper bound of 0.001 representing the largest values from our tidal damping simulations. We then calculated the range of tidal heat fluxes from Equation 4 of Jackson et al. (2008) for both planets and plot the resulting distribution in Figure 14. We find that for and combinations that yield a tidal circularization timescale larger than 2 Gyr and are therefore consistent with the larger eccentricity values preferred for planet b, the upper end of the distribution for tidal heat flux per unit mass is that of Io (TOI-1266 b: W/kg, Io: W/kg, Veeder et al., 2012), while the tidal heat flux in W/m2 at the surface of TOI-1266 b is nearly equal to the insolation flux (Figure 15). This suggests that tidal heating is likely to be important for interpreting the internal and atmospheric composition of TOI-1266 b, with potential implications for long-term atmospheric evolution as well.
@@ 8.1 Bulk Density Constraints
Our updated planet masses and radii do not change the fundamental interpretations of the planet compositions based on the bulk density constraints from C24. The bulk density of TOI-1266 b is likely too low to be explained solely by rock-water mixtures, given the expected upper limit on water mass fraction (WMF) of 50% beyond the snow line for solar composition nebulae (Lodders, 2003), although the expected upper limit on WMF for sub-Neptune exoplanets is still debated (e.g., Bitsch et al., 2021; Burn et al., 2024; Rogers et al., 2025). This implies that TOI-1266 b’s atmosphere likely contains some hydrogen and helium. TOI-1266 c’s smaller radius and larger mass relative to the values reported in C24 now make it easier to match with water-rich scenarios, although it is also consistent with a modest hydrogen-rich envelope (see §8.2). Our updated mass constraint for TOI-1266 c increases the bulk density from 1.6 0.6 g/cm3 to 2.0 0.6 g/cm3, reducing the water-to-rock ratio required to explain the observed mass and radius to 30%. Part of the shift to smaller radii and larger masses for TOI-1266 b and c is also explained by our choice to adopt the stellar parameters from S20, which are based on effective temperature and surface gravity constraints from spectroscopic observations, while C24 adopted stellar parameters from the TIC v8.2 catalog. S20’s stellar mass is larger and their stellar radius is smaller than the values reported in C24.
@@ 8.3 Possible Explanations for the Inverted Architecture of TOI-1266
This picture is additionally complicated by the discovery that TOI-1266 b has a significant tidal heat flux (see §7.4). There are very few compact multi-planet systems with eccentrities as large as our preferred value for TOI-1266 b (e.g., Van Eylen et al., 2019; He et al., 2020). TOI-1266 b’s large tidal heat flux means that its atmosphere may be less massive and/or more metal-rich than implied by the models described in §8, and might help to reduce the disparity in inferred envelope masses relative to TOI-1266 c.
@@ 9.1 TTV Follow-up
There is still much to be gained by additional TTV monitoring of the TOI-1266 system. As shown in Figure 7, we have yet to observe a full TTV super-period for TOI-1266 b. This super-period is caused by its proximity to the 3:1 resonance with TOI-1266 d and has a predicted duration of 2000 days (Lithwick et al., 2012, Equation 5), while the current observational baseline is only 1700 days. This suggests that future TESS and ground-based transit observations over the next decade could significantly improve on current TTV-based mass and eccentricity constraints. To that end, we provide a list of predicted transit times for all three planets from our dynamical fit in Table 5 in the Appendix.
For planets with circular orbits, there is a well-known degeneracy between the planet masses and eccentricities when fitting the observed TTV signal (Lithwick et al., 2012). The nonzero orbital eccentricities of planets b and/or d create a short-period chopping signature in the measured TTVs for planet b, allowing us to break this degeneracy in our TTV fits (e.g., Nesvorný & Vokrouhlický, 2014b; Deck & Agol, 2015).
@@ 9.3 Atmospheric Characterization
If TOI-1266 b has a relatively low atmospheric metallicity (e.g., solar, measured by the abundance of CH4), then the trace abundances of CO and CO2 in the upper atmosphere can be used to constrain the planet’s internal heat flux due to tides (e.g., Fortney et al., 2020; Hu, 2021). If TOI-1266 b has a more metal-rich envelope (e.g. solar), it would have detectable amounts of CO2 even without tidal heating (e.g. Wogan et al., 2024), but the additional tidal heat flux would produce more CO2 and deplete NH3 in the part of the atmosphere probed by transmission spectroscopy. This means that precise abundance measurements for these molecules can be used to constrain the internal heat flux of TOI-1266 b (e.g. Yang & Hu, 2024).
@@ 10 Summary and Conclusions
Our updated planetary interior and composition modeling confirms that TOI-1266 b requires a hydrogen-rich envelope to explain its observed mass and radius. If we assume that this planet has an Earth-like rocky core and a solar composition envelope, this corresponds to a H/He atmospheric mass fraction of %. TOI-1266 c has a modestly higher bulk density ( = 2.24 g/cc, = 1.54 g/cc) and can therefore be matched with either a small hydrogen-rich envelope with an atmospheric mass fraction ( at 95% confidence), or a water-rich envelope with a mass fraction of % from interior structure models, though we caution that much of this range is unsupported by formation models which predict maximum water mass fractions for terrestrial planets . This picture is additionally complicated by our discovery that planet b has a significantly nonzero orbital eccentricity, corresponding to a large tidal heat flux. We simulate the long-term evolution of the system including the effects of tides and find that TOI-1266 b must have relatively weak tidal dissipation (/) in order to maintain this eccentricity over Gyr timescales. In the future, transmission spectroscopy observations of TOI-1266 b with JWST could constrain its internal heat flux, while equivalent observations of TOI 1266 c are needed in order to determine the bulk water content of its envelope. Both planets are among the most favorable temperate sub-Neptune candidates for transmission spectroscopy with JWST, and a detection of atmospheric absorption for either planet would significantly improve our understanding of the TOI-1266 system.
@@ Appendix A Transit Times and Posterior Probability Distributions
TOI-1266 b | D20 | 22 | 1930.692530 | 0.001270000 |
TOI-1266 b | D20 | 23 | 1941.590140 | 0.001480000 |
TOI-1266 b | D20 | 25 | 1963.377320 | 0.002670000 |
TOI-1266 b | MuSCAT | 57 | 2312.013329 | 0.000771316 |
TOI-1266 b | MuSCAT3 | 58 | 2322.907835 | 0.000357439 |
@@
TOI-1266 b | TESS EM | 86 | 2627.962889 | 0.001367946 |
TOI-1266 b | WIRC | 87 | 2638.857341 | 0.000514139 |
TOI-1266 b | MuSCAT3 | 87 | 2638.859811 | 0.001039443 |
TOI-1266 b | TESS EM | 88 | 2649.750889 | 0.001525124 |
TOI-1266 b | TESS EM | 89 | 2660.645312 | 0.001736380 |
@@
TOI-1266 b | TESS SEM | 152 | 3347.007129 | 0.002401229 |
TOI-1266 b | TESS SEM | 154 | 3368.796296 | 0.001593378 |
TOI-1266 b | TESS SEM | 155 | 3379.692673 | 0.002864712 |
TOI-1266 b | TESS SEM | 156 | 3390.588036 | 0.002639571 |
TOI-1266 c | TESS PM | 0 | 1689.960319 | 0.042524505 |
@@
TOI-1266 c | TESS EM | 51 | 2648.845373 | 0.003014198 |
TOI-1266 c | WIRC | 56 | 2742.873075 | 0.015162918 |
TOI-1266 c | TESS SEM | 89 | 3363.304827 | 0.003415930 |
TOI-1266 c | TESS SEM | 90 | 3382.112135 | 0.002607625 |
@@
Figures 17 and 18 show corner plots of the posterior distributions from our joint fit to the TTV and RV data with a 3-planet model. The parameters in our fits were the planet-to-star mass ratio and , , but we have converted these distributions into units of Earth masses and eccentricity for ease of reference in Figure 17. Detrended light curves for all new ground-based transit observations presented in this work, and our stacked TESS transit profiles, are available in the TeX source code documents for this paper.